# Adversarial control of synchronization in complex oscillator networks

### v1

@@ Abstract
This study investigates adversarial attacks, a concept from deep learning, designed to control synchronization dynamics through strategically crafted minimal perturbations. We propose a gradient-based optimization method that identifies small phase perturbations to dramatically enhance or suppress collective synchronization in Kuramoto oscillator networks. Our approach formulates synchronization control as an adversarial optimization problem, computing gradients of the order parameter with respect to oscillator phases to determine optimal perturbation directions. Results demonstrate that extremely small phase perturbations applied to network oscillators can achieve significant synchronization control across diverse network architectures. Our analysis reveals that synchronization enhancement is achievable across various network sizes, while synchronization suppression becomes particularly effective in larger networks, with effectiveness scaling favorably with network size. The method is systematically validated on canonical model networks including scale-free and small-world topologies, and real-world networks representing power grids and brain connectivity patterns. This adversarial framework represents a novel paradigm for synchronization management by introducing deep learning concepts to networked dynamical systems.
Synchronization emerges naturally in many networked systems, where individual components coordinate their behavior to act collectively. Controlling such synchronization in complex networks has been challenging, typically requiring substantial interventions. This study introduces a novel approach inspired by adversarial attacks from artificial intelligence, where tiny, strategically designed perturbations can dramatically alter system behavior. We applied this concept to networks of oscillating systems and discovered that extremely small adjustments can either enhance or suppress synchronization across the entire network. These minimal interventions become particularly effective in larger networks, suggesting excellent scalability for real-world applications. We validated this approach on various network types including power grids and brain networks, demonstrating potential applications for infrastructure stability and managing pathological brain activity. This research offers a new paradigm for controlling complex networked systems with minimal intervention.
@@ I Introduction
External intervention strategies control synchronization without permanently altering system characteristics. Pinning control applies targeted signals to strategically selected nodes to drive network-wide synchronization Wang and Chen (2002); Sorrentino et al. (2007); Xiang et al. (2007); Yang et al. (2010); Lin et al. (2016), with performance depending on appropriate node selection in heterogeneous networks. Time-delayed feedback and global feedback mechanisms (such as mean-field feedback) can promote or suppress synchronization by monitoring the collective state of the system and adjusting control signals accordingly Yeung and Strogatz (1999); Rosenblum and Pikovsky (2004); Taher, Olmi, and Schöll (2019); Franci et al. (2012); Ozawa and Kori (2021), requiring optimization of delay times and feedback strengths, with performance being sensitive to these parameter settings. Noise-induced synchronization exploits the principle that appropriately tuned random perturbations can enhance collective behavior through mechanisms such as stochastic resonance and stochastic synchronization Kawamura, Nakao, and Kuramoto (2007); Esfahani, Shahbazi, and Samani (2012); Lai and Porter (2013); Tyloo, Delabays, and Jacquod (2019), though the stochastic nature of this approach introduces inherent limitations for deterministic control applications.
@@
While these approaches have made valuable contributions, the above challenges remain. To address these limitations, we explore an alternative strategy inspired by adversarial attacks from deep learning. Adversarial attacks involve strategically crafted small perturbations added to inputs that can dramatically alter the outputs of deep neural network models, exploiting inherent system vulnerabilities Yuan et al. (2019); Hirano and Takemoto (2020); Hirano, Koga, and Takemoto (2020); Hirano, Minagi, and Takemoto (2021); Chakraborty et al. (2021); Akhtar et al. (2021); Koga and Takemoto (2022). This concept has recently been extended to dynamics on networks in other domains, specifically applied to voter model dynamics and prisoner’s dilemma games on networks, where small strategic perturbations can significantly alter collective behaviors Chiyomaru and Takemoto (2022, 2023); Mizutaka (2023); Takemoto (2024); Ninomiya et al. (2025). We introduce this adversarial framework to synchronization control in Kuramoto oscillator networks, where the goal is to identify minimal perturbations to the networked dynamics that can disrupt or enhance synchronization by exploiting the system’s intrinsic sensitivities. This approach offers unique characteristics: requiring only weak perturbations to phases while leveraging optimization principles to achieve large-scale effects on collective behavior. By treating synchronization control as an adversarial problem, we introduce a novel theoretical framework that complements existing approaches and provides new insights into the fundamental mechanisms of synchronization manipulation.
@@
In this work, we develop and analyze adversarial attack strategies specifically designed for Kuramoto oscillator networks to achieve synchronization control through phase perturbations. We formulate the adversarial attack as an optimization problem that identifies minimal perturbations to oscillator phases that can maximally disrupt or enhance network synchronization. Our approach combines gradient-based optimization techniques with network analysis to systematically exploit the vulnerabilities inherent in the oscillator networks. We demonstrate that synchronization can indeed be effectively controlled through this adversarial framework via extensive numerical simulations on both model networks and real-world networks. Moreover, we investigate and discuss the effects of network size, average node degree, and network connectivity patterns on the outcomes of adversarial attacks.
@@ III Adversarial attcks
However, obtaining exact phase values and controlling perturbation strength can be challenging. To avoid these limitations, inspired by the fast gradient sign method Goodfellow, Shlens, and Szegedy (2015) for adversarial attacks on deep neural network models, we consider using only the sign of the gradient. Specifically, we add perturbations to each oscillator’s phase at regular time intervals as follows:
@@ IV.1 Network models and simulation setup
We employ networks consisting of oscillators with an average connectivity , unless stated otherwise. Robustness tests confirm that varying network size () or average degree ( does not alter the qualitative behavior. Each oscillator’s intrinsic frequency follows a standard normal distribution , while initial phases are randomly assigned from a uniform distribution over .
@@ IV.2 Adversarial perturbation effects on synchronization transitions
Figure 1 shows the relationship between the order parameter and coupling strength for different perturbation parameters across three network topologies. The most striking observation is that even small perturbation strengths can significantly alter synchronization transitions regardless of network topology. Positive values (red curves) consistently promote synchronization by shifting the transition to lower coupling strengths, while negative values (blue curves) suppress synchronization, requiring higher coupling strengths to achieve the same level of synchronization across all network types. This demonstrates the remarkable effectiveness of adversarial perturbations in controlling synchronization dynamics with minimal intervention strength.
@@ IV.3 Perturbation strength effects
The results demonstrate a striking sensitivity of synchronization dynamics to perturbation strength. For synchronization enhancement (upper panels), increases sharply with even small positive values, showing that minimal perturbations can yield substantial synchronization promotion. Similarly, for synchronization suppression (lower panels), decreases rapidly as the magnitude of negative increases, indicating that small perturbation strengths are sufficient to significantly disrupt collective dynamics. This sharp dependence on underscores the remarkable efficiency of adversarial attacks in manipulating synchronization with minimal intervention.
@@ V Discussion
Our results confirm that borrowing the adversarial attack concept from the deep learning domain and applying it to network synchronization control represents a significant advance. Unlike traditional synchronization control methods that typically require structural modifications Zhou and Kurths (2006); Chen et al. (2008); Zhang and Xiao (2014); Pinto and Saa (2015); Papadopoulos et al. (2017); Ghorban, Hesaam, and Sarbazi-Azad (2022); Ruangkriengsin and Porter (2023); Song et al. (2024); Menara et al. (2022) such as adding or removing links, rewiring connections, or introducing external driving forces Wang and Chen (2002); Sorrentino et al. (2007); Xiang et al. (2007); Yang et al. (2010); Lin et al. (2016); Yeung and Strogatz (1999); Rosenblum and Pikovsky (2004); Taher, Olmi, and Schöll (2019); Franci et al. (2012); Ozawa and Kori (2021); Kawamura, Nakao, and Kuramoto (2007); Esfahani, Shahbazi, and Samani (2012); Lai and Porter (2013); Tyloo, Delabays, and Jacquod (2019), our adversarial perturbation strategy achieves effective control through minimal interventions. While conventional approaches often necessitate global knowledge of network topology or significant alterations to the system architecture, our method operates by applying small amplitude perturbations () to the phases of network nodes, requiring only the computation of gradients and maintaining the original network structure intact.
@@
Despite these promising results, several limitations warrant acknowledgment. Currently, our approach requires access to all network nodes for perturbation, which may not be realistic in many practical scenarios. Developing sparse attack strategies, potentially using gradient magnitudes as selection criteria, represents an important future direction. Additionally, the optimal balance between perturbation parameter and intervention timing requires careful consideration. While larger perturbations and more frequent interventions naturally enhance control effectiveness, practical applications demand minimal interference with the system. The key challenge is achieving robust synchronization control using the smallest possible perturbations applied as infrequently as possible. A deeper theoretical analysis of the attack mechanism would provide valuable insights into these parameter trade-offs and the fundamental principles underlying the observed synchronization control.

### v3

@@ Abstract
This study investigates perturbation strategies inspired by adversarial attack principles from deep learning, designed to control synchronization dynamics through strategically crafted weak perturbations. We propose a gradient-based optimization method that identifies small phase perturbations to dramatically enhance or suppress collective synchronization in Kuramoto oscillator networks. Our approach formulates synchronization control as an optimization problem, computing gradients of the order parameter with respect to oscillator phases to determine optimal perturbation directions. Results demonstrate that extremely small phase perturbations applied to network oscillators can achieve significant synchronization control across diverse network architectures. Our analysis reveals that synchronization enhancement is achievable across various network sizes, while synchronization suppression becomes particularly effective in larger networks, with effectiveness scaling favorably with network size. The method is systematically validated on canonical model networks including scale-free and small-world topologies, and real-world networks representing power grids and brain connectivity patterns. This adversarial framework represents a novel paradigm for synchronization management by introducing deep learning concepts to networked dynamical systems.
Synchronization emerges naturally in many networked systems, where individual components coordinate their behavior to act collectively. Controlling such synchronization in complex networks has been challenging, typically requiring substantial interventions. This study introduces a novel approach inspired by adversarial attack principles from artificial intelligence, where tiny, strategically designed perturbations can dramatically alter system behavior. We applied this concept to networks of oscillating systems and discovered that extremely small adjustments can either enhance or suppress synchronization across the entire network. These weak interventions become particularly effective in larger networks, suggesting excellent scalability for real-world applications. We validated this approach on various network types including power grids and brain networks, demonstrating potential applications for infrastructure stability and managing pathological brain activity. This research offers a new paradigm for controlling complex networked systems with weak intervention.
@@ I Introduction
External intervention strategies control synchronization without permanently altering system characteristics. Pinning control applies targeted signals to strategically selected nodes to drive network-wide synchronization Wang and Chen (2002); Sorrentino et al. (2007); Xiang et al. (2007); Yang et al. (2010); Lin et al. (2016), with performance depending on appropriate node selection in heterogeneous networks. Time-delayed feedback and global feedback mechanisms (such as mean-field feedback) can promote or suppress synchronization by monitoring the collective state of the system and adjusting control signals accordingly Yeung and Strogatz (1999); Rosenblum and Pikovsky (2004); Taher, Olmi, and Schöll (2019); Franci et al. (2012); Ozawa and Kori (2021), requiring optimization of delay times and feedback strengths, with performance being sensitive to these parameter settings. Noise-induced synchronization exploits the principle that appropriately tuned random perturbations can enhance collective behavior through mechanisms such as stochastic resonance and stochastic synchronization Kawamura, Nakao, and Kuramoto (2007); Esfahani, Shahbazi, and Samani (2012); Lai and Porter (2013); Tyloo, Delabays, and Jacquod (2019), though the stochastic nature of this approach introduces inherent limitations for deterministic control applications. The sensitivity of networked oscillators to strategic perturbations has been further demonstrated in studies of network vulnerabilities, where targeted interventions at individual nodes can produce significant changes in global synchronization behavior Tyloo (2023).
@@
While these approaches have made valuable contributions, the above challenges remain. To address these limitations, we explore an alternative strategy inspired by adversarial attack principles from deep learning. Adversarial attacks involve strategically crafted small perturbations added to inputs that can dramatically alter the outputs of deep neural network models, exploiting inherent system vulnerabilities Yuan et al. (2019); Hirano and Takemoto (2020); Hirano, Koga, and Takemoto (2020); Hirano, Minagi, and Takemoto (2021); Chakraborty et al. (2021); Akhtar et al. (2021); Koga and Takemoto (2022). This concept has recently been extended to dynamics on networks in other domains, specifically applied to voter model dynamics and prisoner’s dilemma games on networks, where small strategic perturbations can significantly alter collective behaviors Chiyomaru and Takemoto (2022, 2023); Mizutaka (2023); Takemoto (2024); Ninomiya et al. (2025). We introduce this adversarial framework to synchronization control in Kuramoto oscillator networks, where the goal is to identify small perturbations to the networked dynamics that can disrupt or enhance synchronization by exploiting the system’s intrinsic sensitivities. This approach offers unique characteristics: requiring only weak perturbations to phases while leveraging optimization principles to achieve large-scale effects on collective behavior. By treating synchronization control as an adversarial problem, we introduce a novel theoretical framework that complements existing approaches and provides new insights into the fundamental mechanisms of synchronization manipulation. While our method draws inspiration from adversarial attacks in deep learning, specifically the principle of using small perturbations to achieve large effects, it is fundamentally a gradient-based control strategy that requires real-time access to system states rather than a machine learning technique. We emphasize that our approach represents a conceptual transfer of the adversarial principle rather than establishing direct mathematical analogies between neural networks and oscillator networks.
@@
In this work, we develop and analyze adversarial attack strategies specifically designed for Kuramoto oscillator networks to achieve synchronization control through phase perturbations. We formulate the adversarial attack as an optimization problem that identifies near-minimal perturbations to oscillator phases that can maximally disrupt or enhance network synchronization. Our approach combines gradient-based optimization techniques with network analysis to systematically exploit the vulnerabilities inherent in the oscillator networks. We demonstrate that synchronization can indeed be effectively controlled through this adversarial framework via extensive numerical simulations on both model networks and real-world networks. Moreover, we investigate and discuss the effects of network size, average node degree, and network connectivity patterns on the outcomes of adversarial attacks.
@@ III Adversarial attcks
However, obtaining exact phase values and controlling perturbation strength can be challenging. To avoid these limitations, following the principle of the fast gradient sign method Goodfellow, Shlens, and Szegedy (2015), we consider using only the sign of the gradient. Specifically, we add perturbations to each oscillator’s phase at regular time intervals as follows:
@@ IV.1 Network models and simulation setup
We employ networks consisting of oscillators with an average degree , unless stated otherwise. Robustness tests confirm that varying network size () or average degree ( does not alter the qualitative behavior. Each oscillator’s intrinsic frequency follows a standard normal distribution , while initial phases are randomly assigned from a uniform distribution over .
@@ IV.2 Adversarial perturbation effects on synchronization transitions
Figure 1 shows the relationship between the order parameter and coupling strength for different perturbation parameters across three network topologies. The most striking observation is that even small perturbation strengths can significantly alter synchronization transitions regardless of network topology. Positive values (red curves) consistently promote synchronization by shifting the transition to lower coupling strengths, while negative values (blue curves) suppress synchronization, requiring higher coupling strengths to achieve the same level of synchronization across all network types. This demonstrates the remarkable effectiveness of adversarial perturbations in controlling synchronization dynamics with weak intervention strength.
@@ IV.3 Perturbation strength effects
The results demonstrate a striking sensitivity of synchronization dynamics to perturbation strength. For synchronization enhancement (upper panels), increases sharply with even small positive values, showing that weak perturbations can yield substantial synchronization promotion. Similarly, for synchronization suppression (lower panels), decreases rapidly as the magnitude of negative increases, indicating that small perturbation strengths are sufficient to significantly disrupt collective dynamics. This sharp dependence on underscores the remarkable efficiency of adversarial attacks in manipulating synchronization with weak intervention.
@@ V Discussion
Our results confirm that borrowing the adversarial attack concept from the deep learning domain and applying it to network synchronization control represents a significant advance. Unlike traditional synchronization control methods that typically require structural modifications Zhou and Kurths (2006); Chen et al. (2008); Zhang and Xiao (2014); Pinto and Saa (2015); Papadopoulos et al. (2017); Ghorban, Hesaam, and Sarbazi-Azad (2022); Ruangkriengsin and Porter (2023); Song et al. (2024); Menara et al. (2022) such as adding or removing links, rewiring connections, or introducing external driving forces Wang and Chen (2002); Sorrentino et al. (2007); Xiang et al. (2007); Yang et al. (2010); Lin et al. (2016); Yeung and Strogatz (1999); Rosenblum and Pikovsky (2004); Taher, Olmi, and Schöll (2019); Franci et al. (2012); Ozawa and Kori (2021); Kawamura, Nakao, and Kuramoto (2007); Esfahani, Shahbazi, and Samani (2012); Lai and Porter (2013); Tyloo, Delabays, and Jacquod (2019), our adversarial perturbation strategy achieves effective control through weak interventions. While this approach builds on well-established principles of gradient-based feedback and weak intervention strategies, our contribution lies in the specific formulation and application context that leverages the order parameter structure for distributed phase perturbations. While conventional approaches often necessitate global knowledge of network topology or significant alterations to the system architecture, our method operates by applying small amplitude perturbations () to the phases of network nodes, requiring only the computation of gradients and maintaining the original network structure intact.
It is important to distinguish our approach from related work on network perturbations, particularly Byzantine attacks on oscillator networks Tyloo (2023). While both approaches demonstrate the significant impact that strategic perturbations can have on networked oscillators, they differ fundamentally in purpose and methodology. Byzantine attacks involve malicious actors who completely control individual nodes with arbitrary input signals to disrupt or destabilize synchronization for adversarial purposes. In contrast, our gradient-based framework uses the mathematical structure of the order parameter to determine optimal perturbation directions that enhance or suppress synchronization for beneficial control objectives. Rather than arbitrary signal injection, our method employs principled optimization to achieve desired synchronization states while maintaining the integrity of the network structure. This represents a constructive application of strategic perturbation principles, transforming concepts originally associated with network vulnerabilities into tools for beneficial synchronization management.
@@
While our results demonstrate the effectiveness of adversarial perturbations for synchronization control, we acknowledge important practical considerations for real-world implementation. The current approach requires simultaneous access to all oscillator phases and the ability to perturb all nodes, which presents challenges in practical systems such as power grids or brain networks. We view our current results as establishing a theoretical foundation that opens promising avenues for future developments.
Building on this foundation, several exciting research directions emerge: developing sparse perturbation schemes that achieve effective control by targeting only strategically selected nodes, creating partial-access strategies that operate with limited system information, and designing adaptive methods that can function under realistic constraints. These developments represent natural extensions of our framework that could enable practical synchronization control in real-world networked systems while maintaining the core advantages of weak intervention and distributed control.
Additionally, the optimal balance between perturbation parameter and intervention timing requires careful consideration. In this study, values were selected empirically based on the trade-off between control effectiveness and intervention frequency, as demonstrated in our parameter sweep analysis (Fig. 2). While larger perturbations and more frequent interventions naturally enhance control effectiveness, practical applications demand minimal interference with the system. The key challenge is achieving robust synchronization control using the smallest possible perturbations applied as infrequently as possible. A deeper theoretical analysis of the attack mechanism would provide valuable insights into these parameter trade-offs and the fundamental principles underlying the observed synchronization control.
@@ Acknowledgements.
- Tyloo (2023) M. Tyloo, “Assessing the impact of byzantine attacks on coupled phase oscillators,” Journal of Physics: Complexity 4, 045005 (2023).